# Adversarial Attacks on Reinforcement Learning Agents for Command and Control

### v1

@@ 6 Discussions and Future Work
This directly leads us to two avenues for future work. The first is developing explainable and controllable approaches to COA generation using RL training and developing robustness mechanisms for RL agents that work both during inference and training time. Such mechanisms can be a combination of adversarial perturbation detection and prevention mechanisms that can be deployed on top of pre-existing agents. The second avenue for future work is existing training algorithms can be augmented to train agents that are certifiably robust to malicious noise. In this vein, adversarial training offers a promising alternative to train robust agents. In this work, while we present a preliminary study relating the flatness of the loss landscape of the policy network to it’s apparent robustness, further research is required to quantify the susceptibility of a policy network to noise.

### v2

@@ 6 Discussions and Future Work
This directly leads us to two avenues for future work. The first is developing explainable and controllable approaches to COA generation using RL training and developing robustness mechanisms for RL agents that work both during inference and training time. Such mechanisms can be a combination of adversarial perturbation detection and prevention mechanisms that can be deployed on top of pre-existing agents. The second avenue for future work is existing training algorithms can be augmented to train agents that are certifiably robust to malicious noise. In this vein, adversarial training offers a promising alternative to train robust agents. In this work, while we present a preliminary study relating the flatness of the loss landscape of the policy network to it’s apparent robustness, further research is required to quantify its susceptibility to noise.
@@ 7 Conclusions
{dci}
The Authors declare that there is no conflict of interest.